# On $\operatorname{CAT}(\kappa)$ surfaces

### v2

@@ 5.3. The third proof
Let $\mathcal{E}_{n}$ be the edge graph of $\mathcal{T}_{n}$. According to 2.5, the induced length metric $\widetilde{d}_{n}$ on $\mathcal{E}_{n}$ satisfies
$d(x,y)\leq\widetilde{d}_{n}(x,y)\leq d(x,y)+\varepsilon_{n}$
@@
for all $x,y\in\mathcal{E}_{n}$. Define the surface $X_{n}$ by, for each $T\in\mathcal{T}_{n}$, gluing in a model triangle of constant curvature $\kappa$ having the same edge lengths into the edge graph $\mathcal{E}_{n}$. Identify $\mathcal{E}_{n}$ with a subset of $X_{n}$ in the natural way, and $X_{n}$ with $X$ by choosing some bijection that is the identity map on the edge graph $\mathcal{E}_{n}$. Having made these identifications, denote the new metric on $X$ by $\bar{d}_{n}$.
@@
It follows from the $\operatorname{CAT}(\kappa)$ condition that $d(x,y)\leq\bar{d}_{n}(x,y)$ for all $x,y\in\partial T$ for some triangle $T\in\mathcal{T}_{n}$. Given two points $x,y\in X$, let $\bar{\Gamma}_{n}$ be a curve from $x$ to $y$ whose $\bar{d}_{n}$-length satisfies $\ell_{\bar{d}_{n}}(\Gamma_{n})\leq\bar{d}_{n}(x,y)+\varepsilon_{n}$. We assume that the restriction of $\bar{\Gamma}_{n}$ to each triangle $T\in\mathcal{T}_{n}$ is a $\bar{d}_{n}$-geodesic. Define the curve $\Gamma$ by replacing the subcurve of $\bar{\Gamma}_{n}$ intersecting a given triangle $T$ with the $d$-geodesic connecting the same endpoints. Then $\Gamma$ satisfies $\ell_{d}(\Gamma)\leq\ell_{\bar{d}_{n}}(\bar{\Gamma}_{n})+2\varepsilon_{n}$. Combining these inequalities gives
@@
We conclude that the spaces $X_{n}$ converge uniformly to $X$.
Finally, we claim that the surfaces $X_{n}$ have locally uniformly bounded curvature. Let $x\in X$. Let $V$ be the interior of the union of all triangles in $\mathcal{T}_{1}$ containing $x$, observing that $x\in V$ and that its closure $\overline{V}$ is a closed polygon containing $x$. For all $n\in\mathbb{N}$, let $V_{n}$ denote the set $V$ considered with the metric $\bar{d}_{n}$. Note that each vertex of $X_{n}$ has total angle at least $2\pi$ and hence negative curvature. So all the positive curvature on $X_{n}$ comes from the interior of the faces. Since $\mathcal{T}_{n}$ is a vertex-edge triangulation, we have the uniform bound $|V_{n}|\leq|V_{1}|$. We conclude the positive part of the curvature on $V_{n}$ is at most $\kappa|V_{1}|$. In summary, $X$ is the uniform limit of surfaces $X_{n}$ of locally uniformly bounded curvature, and hence $X$ itself has bounded curvature.
6. Excess-curvature inequality
In this section, we give a proof of 1.2 relating the excess and total curvature of an arbitrary $\operatorname{CAT}(\kappa)$ triangle. First we verify 1.3, that the Hausdorff area of a surface $X$ of bounded curvature without cusp points coincides with the Alexandrov area. This is an application of Reshetnyak’s theorem that the metric tangent at each point is a cone over a circle. Recall that, for a Borel set $A\subset X$, we use $|A|$ to denote the Hausdorff area (i.e., $2$-measure) and $\text{Area}(A)$ to denote Alexandrov area.
Proof of 1.3.
We must show that $|A|=\text{Area}(A)$ for an arbitrary Borel set $A$. Since both Hausdorff area and Alexandrov area are determined by their values on polygons, without loss of generality we may assume that $A$ is a polygon. Let $A\subset X$ be a polygon. We can consider $A$ as its own metric space with the induced length metric; this does not affect either notion of area. Let $\varepsilon>0$ be given. For each $x\in X$, let $B_{x}$ be an open metric ball at $x$ with the property of being $(1+\varepsilon)$-bi-Lipschitz equivalent to a ball in a Euclidean cone over a circle, and let $\mathcal{U}=\{B_{x}:x\in A\}$. Let $\mathcal{T}$ be a triangulation of $A$ subordinate to the cover $\mathcal{U}$ and having mesh $\varepsilon$. Define a polyhedral surface $A_{\varepsilon}$ by replacing each triangle $T\in\mathcal{T}$ with the corresponding model triangle $\bar{T}_{0}$. It is clear that $|A_{\varepsilon}|=\text{Area}(A_{\varepsilon})$. Moreover, we have that each triangle $T\in\mathcal{T}$ is $(1+\varepsilon)$-bi-Lipschitz equivalent to the triangle $\bar{T}_{0}$, which implies that
We now proceed with the proof of 1.2.
Proof of 1.2.
Let $T$ be a triangle with angles $\alpha,\beta,\gamma$. By subdividing if necessary, we may assume by 3.5 that $T$ is convex. In particular, we can now consider $T$ as its own metric space. Let $T_{n}$ be a sequence of model polyhedral approximations converging to $T$ as constructed in the third proof of 1.1; such a sequence can also be found by invoking the general theory of surfaces of bounded curvature. Note that $T_{n}$ is also a triangle, since the edges of $T$ are still geodesics in the polyhedral approximation. Let $\alpha_{n},\beta_{n},\gamma_{n}$ denote the angles in $T_{n}$ corresponding to the same edges in $T$ as $\alpha,\beta,\gamma$, respectively. The approximations $T_{n}$ come with a Lipschitz map $\varphi_{n}\colon T_{n}\to T$. This implies that $\alpha\leq\alpha_{n}$, $\beta\leq\beta_{n}$, $\gamma\leq\gamma_{n}$, and thus that $\delta(T)\leq\delta(T_{n})$.
The rotation of the boundary $\partial T_{n}$ is
$\tau(T_{n})=(\pi-\alpha_{n})+(\pi-\beta_{n})+(\pi-\gamma_{n})+\tau_{0}(T_{n})=2\pi-\delta(T_{n})+\tau_{0}(T_{n}),$
where $\tau_{0}(T_{n})$ is the contribution to the rotation of $\partial T_{n}$ from the interior points of the edges. The fact that each edge is a geodesic implies that $\tau_{0}(T_{n})\leq 0$. In particular, $\delta(T_{n})\leq 2\pi-\tau(T_{n})$. The Gauss–Bonnet theorem for surfaces of bounded curvature states that $\tau(T_{n})+\omega(T_{n})=2\pi$, where $\omega(T_{n})$ is the interior curvature of $T_{n}$. This relation gives
$\delta(T_{n})\leq 2\pi-\tau(T_{n})=\omega(T_{n}).$
Now, the positive part $\omega_{+}(T_{n})$ satisfies $\omega_{+}(T_{n})=\kappa|T_{n}|$. Since $\omega(T_{n})\leq\omega_{+}(T_{n})$ , we conclude that
$\delta(T)\leq\delta(T_{n})\leq\kappa|T_{n}|.$
Finally, the sequence $|T_{n}|$ converges to $|T|$ as $n\to\infty$. This completes the proof. ∎
@@ 7. Approximation by smooth Riemannian surfaces
In this section, we prove that every $\operatorname{CAT}(\kappa)$ surface, $\kappa\in\mathbb{R}$, is the uniform limit of smooth Riemannian surfaces with Gaussian curvature bounded above by $\kappa$ and having locally uniformly bounded area. In Section 5.3, we showed that such a surface is the uniform limit of $\operatorname{CAT}(\kappa)$ model polyhedral surfaces. What remains is to show that one can smoothen out the vertices while retaining the $\operatorname{CAT}(\kappa)$ condition.
@@
Consider the Euclidean ball $B_{\operatorname{Euc}}(\mathbf{0},R)\subset\mathbb{R}^{2}$ for some $R>0$, equipped with the metric $ds^{2}=\lambda(r)(dr^{2}+r^{2}d\theta^{2})$, with $\lambda$ given by (7.1), with $\alpha>1$ and $\kappa\geq 0$. For all $\delta\in(0,R)$ sufficiently small, there is a smooth Riemannian metric $ds_{\delta}^{2}=\lambda_{\delta}(r)(dr^{2}+r^{2}d\theta^{2})$ that agrees with $ds^{2}$ on $B_{\operatorname{Euc}}(\mathbf{0},R)$ outside the Euclidean ball $B_{\operatorname{Euc}}(\mathbf{0},\delta)$ and has Gaussian curvature at most $\kappa$. Moreover, $\lambda_{\delta}$ is an increasing function on $B_{\operatorname{Euc}}(\mathbf{0},\delta)$.
@@
[…]
We claim that this choice of $\lambda_{\delta}$ gives the required properties. It is easy to see that $\lambda_{\delta}(r)=\lambda(r)$ whenever $r\geq\delta$. Moreover, $g_{\delta}(r)=0$ if $r\leq\delta/2$, from which it follows that $\lambda_{\delta}$ is constant on the interval $[0,\delta]$. From this, we see that $ds_{\delta}$ is locally isometric to the plane on $B(\mathbf{0},\delta)$. We conclude that $ds_{\delta}$ is a smooth Riemannian metric. Observe now that
@@
(7.3) $-g_{\delta}^{\prime}(r)-\frac{g_{\delta}(r)}{r}\leq 2\kappa\cdot\lambda_{\delta}(r),$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
We now proceed with the proof of 1.4.
Proof of 1.4.
Let $X$ be a $\operatorname{CAT}(\kappa)$ surface, where $\kappa\geq 0$. Let $(\varepsilon_{n})$ be a sequence of positive reals limiting to 00, and $X_{n}$ the approximating model polyhedral surface given in Section 5.3 for the value $\varepsilon_{n}$. In particular, there is a homeomorphic $\varepsilon_{n}$-isometry $\psi_{n}\colon X_{n}\to X$ with the property that $|\psi_{n}^{-1}(A)|$ converges to $|A|$ for all sets $A\subset X$. For each vertex $v_{i}$ of $X_{n}$, $i\in\mathbb{N}$, let $r_{i}>0$ be such that $B(v_{i},2r_{i})$ does not contain any other vertex of $X_{n}$. Note that this implies that the balls $B(v_{i},r_{i})$ are disjoint. We also require that $\sum_{i=1}^{\infty}r_{i}<\varepsilon_{n}$. The ball $B(v_{i},r_{i})$ is locally isometric to $B_{\operatorname{Euc}}(\mathbf{0},R_{i})$ for some $R_{i}>0$, equipped with the metric given in 7.1. Choose $\delta_{i}>0$ sufficiently small so that $B_{\operatorname{Euc}}(\mathbf{0},\delta_{i})$ has diameter and area at most $\varepsilon_{n}2^{-i}$, which can be done since $\lambda_{\delta}$ is increasing for $r\in(0,\delta)$ and $\lambda(\delta)\to 0$ as $\delta\to 0$. Let $Y_{n}$ be the smooth Riemannian surface obtained by applying 7.1 to each neighborhood $B_{\operatorname{Euc}}(\mathbf{0},\delta_{i})$. Let $V_{i}$ denote the neighborhood in $X_{n}$ on which the metric is altered.
@@
The ideas of 7.1 also apply to smoothing vertices of model polyhedral surfaces with curvature bounded below. Assume that $X$ is a model polyhedral surface with curvature bounded below for some $\kappa\leq 0$. Then the metric is given in a neighborhood of a vertex by (7.1), where $\alpha<1$. Thus, for sufficiently small $r$, one has $\frac{\partial}{\partial r}\log(\lambda(r))>0$ by (7.2). Since $\kappa\leq 0$, the chain of inequalities at the end of the first part of the proof of 7.1 establishes (7.3) with $\leq$” replaced by $\geq$”. We conclude that the Riemannian approximation in 7.1 has curvature bounded below by $\kappa$.

### v3

@@ 6. Excess-curvature inequality
Let $\mathcal{E}_{n}$ be the edge graph of $\mathcal{T}_{n}$. Observe that $d(x,y)\leq\widetilde{d}_{n}(x,y)$ and that, by ˜2.4, the induced length metric $\widetilde{d}_{n}$ on $\mathcal{E}_{n}$ satisfies
$\widetilde{d}_{n}(x,y)\leq d(x,y)+\varepsilon_{n}$
@@
for all $x,y\in\mathcal{E}_{n}$. Define the surface $X_{n}$ by, for each $S\in\mathcal{T}_{n}$, gluing in a model triangle of constant curvature $\kappa$ having the same edge lengths into the edge graph $\mathcal{E}_{n}$. Identify $\mathcal{E}_{n}$ with a subset of $X_{n}$ in the natural way, and $X_{n}$ with $X$ by choosing some bijection that is the identity map on the edge graph $\mathcal{E}_{n}$. Having made these identifications, denote the new metric on $X$ by $\bar{d}_{n}$. Let $S_{n}^{j}$, $1\leq j\leq k_{n}$, denote the collection of faces of $T_{n}$, each of which is a triangle of constant curvature $\kappa$.
@@
It follows from the $\operatorname{CAT}(\kappa)$ condition that $d(x,y)\leq\bar{d}_{n}(x,y)$ for all $x,y\in\partial S$ for each triangle $S\in\mathcal{T}_{n}$. Given two points $x,y\in X$, let $\bar{\Gamma}_{n}$ be a curve from $x$ to $y$ whose $\bar{d}_{n}$-length satisfies $\ell_{\bar{d}_{n}}(\Gamma_{n})\leq\bar{d}_{n}(x,y)+\varepsilon_{n}$. We assume that the restriction of $\bar{\Gamma}_{n}$ to each triangle $S\in\mathcal{T}_{n}$ is a $\bar{d}_{n}$-geodesic. Define the curve $\Gamma$ by replacing the subcurve of $\bar{\Gamma}_{n}$ intersecting a given triangle $T$ with the $d$-geodesic connecting the same endpoints. Then $\Gamma$ satisfies $\ell_{d}(\Gamma)\leq\ell_{\bar{d}_{n}}(\bar{\Gamma}_{n})+2\varepsilon_{n}$. Combining these inequalities gives
@@
We conclude that the sequence of surfaces $T_{n}$ converges uniformly to $T$ with its induced length metric.
Note that $T_{n}$ is also a triangle, since the edges of $T$ are still geodesics in the polyhedral approximation. Let $\alpha_{n},\beta_{n},\gamma_{n}$ denote the angles in $T_{n}$ corresponding to the same edges in $T$ as $\alpha,\beta,\gamma$, respectively. The $\operatorname{CAT}(\kappa)$ condition implies that $\alpha\leq\alpha_{n}$, $\beta\leq\beta_{n}$, $\gamma\leq\gamma_{n}$, and thus that $\delta(T)\leq\delta(T_{n})$. The excess moreover satisfies
$\delta(T_{n})\leq\sum_{j=1}^{k_{n}}\delta(S_{n}^{j});$
see Lemma III.9 in [1], noting that the proof applies to our situation.
We observe that $\delta(S_{n}^{j})=\kappa|S_{n}^{j}|$ for each triangle $S_{n}^{j}$ by Girard’s Theorem and its generalization. Combining these inequalities gives
$\delta(T)\leq\delta(T_{n})\leq\sum_{j=1}^{k_{n}}\delta(S_{n}^{j})=\sum_{j=1}^{k_{n}}\kappa|S_{n}^{j}|=\kappa|T_{n}|.$
Letting $n\to\infty$, the right-hand side converges to $|T|$. We conclude that $\delta(T)\leq\kappa|T|$. ∎
@@ 7. Approximation by smooth Riemannian surfaces
In this section, we prove that every $\operatorname{CAT}(\kappa)$ surface, $\kappa\in\mathbb{R}$, is the uniform limit of smooth Riemannian surfaces with Gaussian curvature bounded above by $\kappa$ and having locally uniformly bounded area. As already discussed in the proof of ˜1.2, it is well-known that such a surface is the uniform limit of $\operatorname{CAT}(\kappa)$ model polyhedral surfaces. What remains is to show that one can smoothen out the vertices while retaining the $\operatorname{CAT}(\kappa)$ condition.
@@
Consider the Euclidean ball $B_{\operatorname{Euc}}(\mathbf{0},R)\subset\mathbb{R}^{2}$ for some $R>0$, equipped with the metric $ds^{2}=\lambda(r)(dr^{2}+r^{2}d\theta^{2})$, with $\lambda$ given by (7.1), with $\alpha>1$. For all $\delta\in(0,R)$ sufficiently small, there is a smooth Riemannian metric $ds_{\delta}^{2}=\lambda_{\delta}(r)(dr^{2}+r^{2}d\theta^{2})$ that agrees with $ds^{2}$ on $B_{\operatorname{Euc}}(\mathbf{0},R)$ outside the Euclidean ball $B_{\operatorname{Euc}}(\mathbf{0},\delta)$ and has Gaussian curvature at most $\kappa$. Moreover, $\lambda_{\delta}$ is an increasing function on $B_{\operatorname{Euc}}(\mathbf{0},\delta)$.
@@
$\lambda_{\delta}(r)=\lambda(\delta)\exp\left(\int_{\delta}^{r}g_{\delta}(t)\,dt\right).$
We claim that this choice of $\lambda_{\delta}$ gives the required properties. It is easy to see that $\lambda_{\delta}(r)=\lambda(r)$ whenever $r\geq\delta$. Moreover, $g_{\delta}(r)=0$ if $r\leq\delta/2$, from which it follows that $\lambda_{\delta}$ is constant on the interval $0,\delta]$. From this, we see that $ds_{\delta}$ is locally isometric to the plane on $B(\mathbf{0},\delta)$. We conclude that $ds_{\delta}$ is a smooth Riemannian metric. Observe now that
@@
(7.3) $g_{\delta}^{\prime}(r)-\frac{g_{\delta}(r)}{r}\leq 2\kappa\cdot\lambda_{\delta}(r),$
@@
$\left(\frac{\partial}{\partial r}\log(\lambda(r))\right)\varphi_{\delta}^{\prime}(r)-\left(\frac{\partial^{2}}{\partial r^{2}}\log(\lambda(r))\right)\varphi_{\delta}(r)-\frac{\left(\frac{\partial}{\partial r}\log(\lambda(r))\right)\varphi_{\delta}(r)}{r}\leq 2\kappa\cdot\lambda(\delta)\exp\left(\int_{\delta}^{r}g_{\delta}(t)\,dt\right).$
@@
(7.4) $\lambda(r)\varphi_{\delta}(r)\leq\lambda(\delta)\exp\left(\int_{\delta}^{r}g_{\delta}(t)\,dt\right).$
@@
$\exp\left((\log\lambda(r))\varphi_{\delta}(r)\right)$ $=\exp\left(\left(\log\lambda(\delta)+\int_{\delta}^{r}\frac{\partial}{\partial t}\log\lambda(t)\,dt\right)\varphi_{\delta}(r)\right)$
$\leq\exp\left(\left(\log\lambda(\delta)+\int_{\delta}^{r}\frac{\partial}{\partial t}\log\lambda(t)\varphi_{\delta}(t)\,dt\right)\right)$
$=\lambda(\delta)\exp\left(\int_{\delta}^{r}g_{\delta}(t)\,dt\right).$
@@
Keep the same definition for $\lambda_{\delta}$, except that we use the new formula for $g_{\delta}$. The conformal weight $\lambda_{\delta}$ on $B_{\text{Euc}}(\mathbf{0},\delta/2)$ is a rescaling of the hyperbolic metric by the factor $\lambda(\delta)\exp\left(\int_{\delta}^{\delta/2}g_{\delta}(t)\,dt\right)\leq\lambda(\delta/2)$, which is less than $1$ since $\delta<1$. It follows that $K_{\delta}(r)\leq\kappa$ whenever $r\leq\delta/2$.
@@
$\exp\left(\left(\log\lambda(\delta)+\int_{\delta}^{r}\frac{\partial}{\partial t}\log(\lambda(t))\,dt\right)\varphi_{\delta}(r)+\left(\log\bar{\lambda}(\delta)+\int_{\delta}^{r}\frac{\partial}{\partial t}\log(\bar{\lambda}(t))\,dt\right)(1-\varphi_{\delta}(r))\right).$
@@
$\exp\left(\log\lambda(\delta)+\left(\int_{\delta}^{r}\frac{\partial}{\partial t}\log(\lambda(t))\varphi_{\delta}(t)\,dt\right)+\left(\int_{\delta}^{r}\frac{\partial}{\partial t}\log(\bar{\lambda}(t))(1-\varphi_{\delta}(t))\,dt\right)\right)$
$=$ $\lambda(\delta)\exp\left(\int_{\delta}^{r}g_{\delta}(t)\,dt\right).$
@@
We now proceed with the proof of ˜1.4.
Proof of ˜1.4.
Let $X$ be a $\operatorname{CAT}(\kappa)$ surface, where $\kappa\geq 0$. Let $(\varepsilon_{n})$ be a sequence of positive reals limiting to 0, and $X_{n}$ the approximating model polyhedral surface for the value $\varepsilon_{n}$. In particular, there is a homeomorphic $\varepsilon_{n}$-isometry $\psi_{n}\colon X_{n}\to X$ with the property that $|\psi_{n}^{-1}(A)|$ converges to $|A|$ for all sets $A\subset X$. For each vertex $v_{i}$ of $X_{n}$, $i\in\mathbb{N}$, let $r_{i}>0$ be such that $B(v_{i},2r_{i})$ does not contain any other vertex of $X_{n}$. Note that this implies that the balls $B(v_{i},r_{i})$ are disjoint. We also require that $\sum_{i=1}^{\infty}r_{i}<\varepsilon_{n}$. The ball $B(v_{i},r_{i})$ is locally isometric to $B_{\operatorname{Euc}}(\mathbf{0},R_{i})$ for some $R_{i}>0$, equipped with the metric given in ˜7.1. Choose $\delta_{i}>0$ sufficiently small so that $B_{\operatorname{Euc}}(\mathbf{0},\delta_{i})$ has diameter and area at most $\varepsilon_{n}2^{-i}$, which can be done since $\lambda_{\delta}$ is increasing for $r\in(0,\delta)$ and $\lambda(\delta)\to 0$ as $\delta\to 0$. Let $Y_{n}$ be the smooth Riemannian surface obtained by applying ˜7.1 to each neighborhood $B_{\operatorname{Euc}}(\mathbf{0},\delta_{i})$. Let $V_{i}$ denote the neighborhood in $X_{n}$ on which the metric is altered.
@@
The ideas of ˜7.1 also apply to smoothing vertices of model polyhedral surfaces with curvature bounded below. Assume that $X$ is a model polyhedral surface with curvature bounded below for some $\kappa\leq 0$. Then the metric is given in a neighborhood of a vertex by (7.1), where $\alpha<1$. Thus, for sufficiently small $r$, one has $\frac{\partial}{\partial r}\log(\lambda(r))>0$ by (7.2). Since $\kappa\leq 0$, the chain of inequalities at the end of the first part of the proof of ˜7.1 establishes (7.3) with $\leq$” replaced by $\geq$”. We conclude that the Riemannian approximation in ˜7.1 has curvature bounded below by $\kappa$.